# Dynamic Order Template Prediction for Generative Aspect-Based Sentiment Analysis

### v1

@@ Abstract
Aspect-based sentiment analysis (ABSA) assesses sentiments towards specific aspects within texts, resulting in detailed sentiment tuples. Previous ABSA models often use static templates to predict all of the elements in the tuples, and these models often fail to accurately capture dependencies between elements. Multi-view prompting method improves the performance of ABSA by predicting tuples with various templates and then ensembling the results. However, this method suffers from inefficiencies and out-of-distribution errors. In this paper, we propose a Dynamic Order Template (DOT) method for ABSA, which dynamically generates necessary views for each instance based on instance-level entropy. Ensuring the diverse and relevant view generation, our proposed method improves F1-scores on ASQP and ACOS datasets while significantly reducing inference time.
@@
[…]
1 Introduction
Aspect-based sentiment analysis (ABSA) aims to identify sentiment for the aspects in a given text rather than simply classifying the overall sentiment of the entire text. ABSA research evolves to generate quadruples consisting of four elements: 1) Aspect ($A$), 2) Category ($C$) for the type of $A$, 3) Opinion ($O$) for $A$, and 4) Sentiment ($S$) for $A$. Many recent studies such as T5-paraphrase (Zhang et al., 2021b) tackle this problem using generative models. These approaches usually get review sentences as input and output the span of quadruples as fixed order paraphrased form, such as "$C$ is $S$ because $A$ is $O$(Zhang et al., 2021a). However, this static single-order template cannot express the dependence between elements as in Figure 1 due to the autoregressive nature of transformer (Vaswani et al., 2017). Moreover, the model’s output can heavily depend on the order of generating each element Hu et al. (2022a).
Figure 1: Comparison of three different generative ABSA methods. 1) static single-view (T5-paraphrase), 2) static multi-view (MvP), and 3) dynamic-view prediction (ours).
Multi-view prompting (Gou et al., 2023) (MvP) deals with this issue by constructing order templates as a channel for "viewing" different perspectives in a sentence. As shown in Figure 1, MvP permutes all possible element orders and sorts them based on the dataset-level entropy of the pre-trained model. Using this entropy, MvP samples top-k orders and adds these orders as a prompt template. During the inference time, MvP conducts majority votes on generated sentiment tuples with various templates. Through this ensemble approach, MvP utilizes the intuition of solving problems from different views in human reasoning and decision Stanovich and West (2000), resulting in significantly higher performance. However, we find that this static multi-view approach of MvP has several drawbacks: 1) Inefficient: Even for samples where the answer can be easily found and multiple views are not necessary, this method generates the same number of views, resulting in unnecessary computation that increases the inference time. 2) Prone to Distribution shift: MvP uses the number of views k as a hyperparameter, applying the same k value across all datasets during training and inference. However, since the optimal number of ensemble models varies according to the data distribution, it requires manual adjustment of the k value for each dataset Shahhosseini et al. (2022), which hinders the transferability to other datasets.
To resolve the aforementioned shortcomings, we propose a Dynamic Order Template (DOT) method for ABSA that combines the advantages of both single-view and multi-view approaches. By prioritizing multiple views based on instance-level entropy, DOT aims to generate only the necessary number of views for each instance during inference. For an example that contains only one tuple as in Figure  1, DOT dynamically creates only one order template (i.e. view) that is necessary for predicting the tuple. After generating the views, DOT generates tuples using the order templates that correspond to the views. This phase operates in a multi-view manner, enabling us to retain the benefits of previous multi-view methods.
Extensive experiments on five widely used sentiment quadruple prediction datasets, derived from ASQP Pontiki et al. (2016); Zhang et al. (2022) and ACOS (Cai et al., 2021, 2023), demonstrate that our method show state-of-the arts performance with significantly lower inference time compared to previous multi-view inference work. Moreover, we show that our method is robust to distribution shift compared to previous methods.
2 Method
Figure 2: Overview of our proposed two stage method. We use two T5 models for each stage: one for generating views, the other for generating sentiment tuples.
Our proposed Dynamic Order Template (DOT) method is composed of two stages as in Figure 2. The first stage predicts the number of order templates to be used as a prompt. (Sec 2.1) Using the order templates, second stage predicts the sentiment tuples. We train the model to generate distinct tuples for each order template, enabling efficient aggregation. (Sec 2.2) For both stages, we map sentiment tuples ($A$, $C$, $S$, $O$) to marker tokens $[A]$, $[C]$, $[S]$, and $[O]$ respectively. For the cases of the instances that contain multiple sentiment tuples, we indicate each tuple with the respective tokens and concatenate the targets with [SSEP] tokens.
2.1 Stage 1: Generating the Number of Views
We assume that the number of sentiment tuples present for each instance corresponds to the required number of views. In other words, we consider that one separate view is necessary for predicting each tuple. We define this individual view as the prediction order for each element of the sentiment tuple as shown in Figure 2. This allows each prediction order to correspond one-to-one with a sentiment tuple in second stage. We set the required number of sentiment tuples $K_{i}$ in the $i^{th}$ instance.
@@ 2.1 Stage 1: Generating the Number of Views
To predict the number of tuples $K_{i}$, we begin by examining the entropy of all possible views generated through permutations. Specifically, we calculate entropy of each view $v$ in instance-level with vanilla T5  Hu et al. (2022a) by calculating conditional generation probability as follows:
$\mathcal{E}_{i,v}$ $=p(\boldsymbol{y_{i,v}}|\boldsymbol{x_{i}}),$ (1)
where $E_{i,v}$ denotes the entropy score of $v$ in the permutation set for $i^{th}$ instance. $y_{i,v}$ is a permuted sentiment tuple in $i^{th}$ instance, based on $v$. At this time, we note that existing ABSA methods have generally struggled to extract $O$ (Chebolu et al., 2023). Based on this observation, we consider that $O$ might hinder the calculation of entropy score. Hence, we calculate entropy scores for $v$ without $O$ (i.e. ($A$, $C$, $S$)) in this stage. We provide a detailed analysis on excluding [O] in Appendix B.2. After computing the entropy, we sort the views by the entropy in ascending order and map each element into marker tokens to get the ranked view $P^{(1)}_{i}$. And then we sample top $K_{i}$ views for each sample $X_{i}$ and concatenate these views as an order template.
@@
Using the original input sentence $X_{i}$, we train the T5 model to generate the first-stage target $y^{(1)}_{i}$. The format of $y^{(1)}_{i}$ is as follows:
@@
where $P^{(1)}_{i,K_{i}}$ denotes the order template for the $j^{th}$ order template of $i^{th}$ instance in the first stage. We set the loss function to train the T5 model as in Equation (2), where $\left\lvert B\right\rvert$ denotes the batch size of the model. The scaling factor is omitted for simplicity.
@@ 2.2 Stage 2: Sentiment Tuple Generation
In the second stage, the model predicts the sentiment tuple of given instance using the number of generated views in the first stage. Different from the first stage, we need to generate all elements in sentiment quadruples including $O$ in this stage. Hence, we re-rank all views to pick $K_{i}$ views including $O$ (i.e. ($A$, $C$, $S$, $O$)). Here, we use the same ranking strategy as in the first stage using the entropy and denote ranked view set as $P^{(2)}_{i}$. We sample top $K_{i}$ views from ranked view and these views are mapped into sentiment tuples one by one, making the model to learn different order template should generate different tuples. We then concatenate the sampled views and add them as a prompt $P_{i}$ to original input sentence $X_{i}$. We design the second stage target $y^{(2)}_{i}$ by placing the corresponding element next to each marker token [SSEP] within $P_{i}$ as follows.
$P_{i}=\begin{aligned} &P^{(2)}_{i,1}\,\text{[SSEP]}\,P^{(2)}_{i,2}\,\text{[SSEP]}\,\ldots\,P^{(2)}_{i,K_{i}},\end{aligned}$
$y^{(2)}_{i}=\begin{aligned} &P^{(2)}_{i,1}\,tuple_{1}\,\text{[SSEP]}\,\ldots\,P^{(2)}_{i,K_{i}}\,tuple_{K_{i}},\end{aligned}$
where $P^{(2)}_{i,K_{i}}$ represents the order template for the $j^{th}$ order template of $i^{th}$ instance in the second stage, and $tuple_{j}$ is the sentiment tuple correspond to $P^{(2)}_{i,K_{i}}$. We design the loss function for training another T5 model in second stage as follows.
@@ 2.3 Two-stage inference
During inference time, two stages are conducted sequentially. In the first stage, we generate the necessary views $\hat{y^{(1)}}$ based on the predicted number of views $\hat{K}$ (i.e. the number of tuples). Given $\hat{K}$, we sample top $\hat{K}$ orders from ranked order set $\hat{P}$ and construct order template $\hat{P}$. Note that we rank orders in full elements for each instance. Lastly, we directly add $\hat{P}$ to inference sentence and generate targeted sentiment tuples in second stage.
@@ 3.1 Benchmark Datasets
We adopt two widely used ABSA datasets: ASQP and ACOS, where the task is to predict sentiment quadruples. For ASQP task, we use rest15(R15) and rest16(R16) datasets released from Pontiki et al. (2016); Zhang et al. (2022). For ACOS task, we use laptop16(Lap) and rest16(Rest) datasets constructed by Cai et al. (2021); Pontiki et al. (2016). Also, we adopt additional ACOS dataset(MR) from MEMD restaurant data Xu et al. (2023) which uses a different source from the previous datasets.
@@ 3.2 Baselines
We compare our method against several strong baselines for ABSA as follows. Paraphrase Zhang et al. (2021a) formulates a novel paraphrase generation process for ABSA with a single fixed order. Seq2Path Mao et al. (2022) generates sentiments tuples as multiple paths of a tree, and automatically selects valid one. DLO Hu et al. (2022b) augments data via the multiple order templates. MvP aggregates sentiment tuples generated from different orders of prompts via ensembling. AugABSA Wang et al. (2023) generates a original text based on augmented sentiment quadruples. Also, we benchmark popular LLMs such as GPT-3.5, LLaMa-3 Team (2024), and Mistral-7b Jiang et al. (2023). Detailed setups for LLMs are described in Appendix D.
@@ 3.3 Implementation Details
We utilize the pre-trained T5-base (Raffel et al., 2020) model as the backbone for the first stage. We also use the model trained in the first stage as the backbone for the second stage, allowing us to leverage a tuned initial point for the ABSA dataset to have the regularization effect inspired by Fu et al. (2023). Also, we eliminate irregularities in tuples through stop-word filtering in the second stage. Please see Appendix A for more details.
@@ 3.4 Results
Methods ASQP ACOS Time(s)
R15 R16 Lap Rest MR
Paraphrase 46.93 57.93 43.51 61.16 57.38 40.63
Seq2Path - - 42.97 58.41 - -
DLO 48.18 59.79 43.64 59.99 57.07 260.74
MvP 51.04 60.39 43.92 61.54 58.12 2161.81
AugABSA 50.01 60.88 - - - -
GPT 3.5-turbo 34.27 36.71 16.00 37.71 - -
LLaMa3 8b 37.52 47.60 40.07 54.06 - -
Mistral 7b 44.14 51.96 39.02 53.02 - -
DOT (Ours) 51.91 61.24 44.92 59.25 58.25 298.17
Table 1: F1 scores for ABSA on five datasets. The best results are in bold and the second best are underlined. We conduct experiments with 5 different seeds and report the average of the outcomes. Time denotes the averaged inference time.
F1 score
We use F1 score, which is a standard metric for ABSA, to measure the performance of the systems. As demonstrated in Table 1, our proposed method outperforms all baselines and achieves state-of-the-art performance across four datasets for ABSA. Our model shows slightly reduced performance on the Rest dataset, which we attribute to the substantial number of implicit aspects and opinions within this dataset. Additionally, the size of the Rest dataset is relatively small, being less than half that of the Lap or MR datasets, which are derived from the same ACOS dataset.
@@
We also measure inference time using T5-base model for all baselines. We check inference time for each dataset, and average them. As in Table 1, we dramatically reduce inference time particularly compared to the multi-view methods such as MvP Gou et al. (2023), by predicting solely the necessary number of views for each sample. On the other hand, in terms of single view inferences Zhang et al. (2021a), we significantly improve the F1 score performance while suppressing the rate of increase in inference time. We also provide more details about computing the inference time in Appendix C.
3.5 Analysis
Ablation study
To further investigate the effectiveness of each component of our framework, we conduct an ablation study and present the average F1 score across the datasets in Table 2. We first unify the two stages into one, directly generating multiple order templates and tuples without including order prompting to validate the effect of stage division. Additionally, we evaluate the results of directly using the generated views from the first stage, omitting the sampling of new order templates. Lastly, we exclude the multi-view approach by training and testing our model using the order template with the lowest entropy for each instance. By observing the gaps between these variants with the original model, we verify the effectiveness of each component of our method.
Model Configuration Average F1
Full Model 53.94
w/o stage division 52.73 (-1.21)
w/o re-sampling 52.47 (-1.47)
w/o multi view 53.31 (-0.63)
Table 2: Ablation study for the proposed method.
Distribution Shift
To examine the effect of distribution shift of each model, we conduct an in-depth experiment on cross-dataset evaluation. We group the datasets into two groups based on their source: SemEval Pontiki et al. (2016) (R15, R16, Rest) and Yelp (MEMD). Then we assess the performance between these groups by training on one group and testing on the other in a zero-shot setting. For the MvP model, we vary the number of views used for ensembling into 3, 9, and 15 (default) to measure the sensitivity of this number in static multi-view methods. Additionally, we evaluate T5-paraphrase which uses a static single order. Table 3 demonstrates that our model significantly outperforms the baselines in cross-dataset evaluation. While T5-paraphrase experiences a smaller performance drop compared to the others, it still lags behind our method. In particular, MvP exhibits significant performance degradation, irrespective of the number of views. From these experiments, we show that our model can effectively find the optimal number of views even for the out-of-domain datasets.
@@ 3.5 Analysis
Table 3: Cross-dataset evaluation results for validating the effect of distribution shift.
Figure 3: Case study for three main types of results. Blue one denotes correct, red one denotes incorrect, and the yellow one denotes irrelevant.
Case Study
The second shows an example predicts require fewer views, but the example actually requires more views. Our analysis reveals that such cases frequently occurs with implicit $O$. As shown in Table 1, this suggests that our model’s performance might lag behind other baselines on the ACOS Rest16 dataset, which contains many samples with implicit $A$ and $O$. Additionally, the model struggles with predicting infrequent $C$ in the training set. Incorporating the concept of self-information and defining the necessary number of views based on the ’amount of information in a sample’ could effectively address this issue.
The final case involves cases with multiple sentiment tuples and longer lengths. We explain that errors in this scenario stem from two main reasons. Firstly, longer sentences include extended phrases that modify $A$ or $O$. Including all these modifiers as elements often leads to errors, a common problem across different models that requires an alternative solution. Secondly, errors occur when the number of tuples is incorrectly predicted in the first stage. If the predicted number of tuples is insufficient, some target sentiment tuples might be overlooked. Conversely, overestimation leads to the extraction of irrelevant aspects, as depicted in the figure. However, we optimize the first stage to reduce tuple count errors, which helped mitigate performance drops by minimizing incorrectly generated or overlooked tuples.

### v2

@@ Abstract
Aspect-based sentiment analysis (ABSA) assesses sentiments towards specific aspects within texts, resulting in detailed sentiment tuples. Previous ABSA models often used static templates to predict all the elements in the tuples, and these models often failed to accurately capture dependencies between elements. Multi-view prompting method improves the performance of ABSA by predicting tuples with various templates and then assembling the results. However, this method suffers from inefficiencies and out-of-distribution errors. In this paper, we propose a Dynamic Order Template (DOT) method for ABSA, which dynamically creates an order template that contains only the necessary views for each instance. Ensuring the diverse and relevant view generation, our proposed method improves F1 scores on ASQP and ACOS datasets while significantly reducing inference time.***Our implementation is publicly available at https://github.com/imsongpasimin/DOT
@@
Yonghyun Jun  and Hwanhee Leethanks: Corresponding author. Department of Artificial Intelligence, Chung-Ang University {zgold5670, hwanheelee}@cau.ac.kr
1 Introduction
Aspect-based sentiment analysis (ABSA) aims to identify the sentiment of aspects in a given text rather than simply classifying the overall sentiment of the entire text. ABSA research evolves to generate quadruples consisting of four elements: 1) Aspect ($A$), 2) Category ($C$) for the type of $A$, 3) Opinion ($O$) for $A$, and 4) Sentiment ($S$) for $A$. Many recent studies, such as T5-paraphrase, tackle this problem using generative models (Zhang et al., 2021b). These approaches usually get review sentences as input and output the span of quadruples in fixed order form, such as "$C$ is $S$ because $A$ is $O$(Zhang et al., 2021a). However, this static single-order template cannot express the dependence between elements as in Figure 1 due to the autoregressive nature of the transformer (Vaswani et al., 2017). Moreover, the model output can heavily depend on the order of generation of each element Hu et al. (2022).
Figure 1: Comparison of three different generative ABSA methods. 1) static single-view, 2) static multi-view, and 3) dynamic-view prediction (ours).
Multi-view prompting (Gou et al., 2023) (MvP) deals with this issue by constructing order templates as a channel for "viewing" different perspectives in a sentence. As shown in Figure 1, MvP permutes all possible element orders and sorts them based on the entropy of the pre-trained model at the dataset level. Using this entropy, MvP samples top-k orders and adds these orders as a prompt template. During inference time, MvP conducts majority votes on generated sentiment tuples with various templates. Through this ensemble approach, MvP uses the intuition of solving problems from multiple views in human reasoning and decision Stanovich and West (2000), resulting in enhanced performance. However, we find that this static multi-view approach of MvP has several drawbacks: 1) Inefficiency: Even for samples where the answer can be easily found and multiple views are not necessary, this method generates the same number of views, resulting in unnecessary computation that increases the inference time. 2) Limited Transferability: MvP uses the number of views k as a hyperparameter, applying the same k value across all datasets during training and inference. However, since the optimal number of ensemble models varies according to the data domain, it requires manual adjustment of the k value for each dataset Shahhosseini et al. (2022), which hinders the transferability to other datasets.
To resolve the aforementioned shortcomings, we propose a Dynamic Order Template (DOT) method for ABSA that combines the advantages of both single-view and multi-view approaches. By prioritizing multiple views based on instance-level entropy, DOT aims to generate only the necessary number of views for each instance during inference. For an example that contains only one tuple, as in Figure 1, DOT dynamically creates only one view as an order template necessary to predict the tuple. After generating the views, DOT generates tuples using the views inside the order template. This phase operates in a multi-view manner, enabling us to retain the benefits of previous multi-view methods. Extensive experiments on four widely used sentiment quadruple prediction datasets, derived from ASQP Pontiki et al. (2016); Zhang et al. (2022), ACOS (Cai et al., 2021a, 2023), and MEMD-ABSA Cai et al. (2023), demonstrate that our method shows state-of-the-art performance with significantly lower inference time compared to the multi-view approach. Moreover, we show that our method is robust to domain shift compared to previous methods, resulting in higher transferability.
2 Method
Figure 2: Overview of our proposed two stage method. We use two T5 models for each stage: one for generating the initial order template, the other for forming the final order template and generating sentiment tuples.
@@ 2 Method
Also, for the instances that contain multiple sentiment tuples, we indicate each tuple with the respective tokens and concatenate the targets with [SSEP] tokens. It is important to note that null or missing values in the datasets are encoded as the literal string "null". During template construction, our method maps this string directly to the designated marker tokens, eliminating the need for any special handling.
2.1 Stage 1: Generating Order Template
We assume that the number of sentiment tuples $K_{i}$ in $i^{th}$ instance present for each instance corresponds to the required number of views. Consequently, we are able to divide the complex sentiment tuple generation task into two relatively simpler subtasks: 1) predicting the number of sentiment tuples in the sentence, and 2) generating sentiment tuples with the exact number of first stage predictions. Building separate models specialized for each subtask and enabling collaboration between them was instrumental in achieving significant performance improvements. In this framework, each view is interpreted as the ordered scaffold for generating a single sentiment tuple, as depicted in Figure 2. This allows each prediction order to correspond one-to-one with a sentiment tuple in the second stage.
We observe that instead of directly using the value of $K_{i}$ as the target, sampling views corresponding to $K_{i}$ as the target for the model to generate leads to more accurate prediction of the number of tuples in the given instance (More details are in Appendix C). To establish the view sampling strategy, we start by ranking all possible views generated through permutations through each entropy score, following  Hu et al. (2022).
@@ 2.1 Stage 1: Generating Order Template
Specifically, we calculate entropy of each view $v$ in instance-level with vanilla T5 by calculating conditional generation probability as follows:
$\mathcal{E}_{i,v}$ $=-\sum P(v|{x_{i}})\log P(v|{x_{i}})$ (1)
Here, $\mathcal{E}_{i,v}$ is the entropy of the total sequence when the $i_{th}$ instance is input into the T5 model and $v$ is the output. At this time, we note that actually utilizing only $A$, $C$, $S$ during the first stage notably facilitates the training process in the second stage. We provide a detailed analysis on excluding $O$ in Appendix C.2. After computing the entropy, we sort the views by the entropy in ascending order to get the ranked set of views $P^{(1)}_{i}$. And then we sample the top $K_{i}$ views for each sample and concatenate these views as an order template.
@@
Using the original $i^{th}$ input sentence, we train the T5 model to generate the first-stage target $y^{(1)}_{i}$ as follows:
@@
where $P^{(1)}_{i,K_{i}}$ denotes $K_{i}^{th}$ view in $P^{(1)}_{i}$. We set the loss function to train the T5 model as in Equation (2), where $\left\lvert B\right\rvert$ denotes the batch size of the model. The scaling factor is omitted for simplicity.
@@ 2.2 Stage 2: Sentiment Tuple Generation
In the second stage, the model is trained to generate the sentiment tuple of a given instance using the number of sentiment tuples (i.e. $K_{i}$). Different from the first stage, we need to generate all elements in sentiment quadruples including $O$ in this stage. Hence, we re-rank all views to pick $K_{i}$ views including $O$ (i.e. ($A$, $C$, $S$, $O$)).
Here, we adopt the same strategy as in the first stage, using entropy to form a ranked set of views, $P^{(2)}_{i}$. We then sample top $K_{i}$ views from $P^{(2)}_{i}$ and add them as an order template prompt $P_{i}$ to original input sentence.
We design the second stage target $y^{(2)}_{i}$ by aligning each sentiment tuple with an order template, ensuring that the model learns to generate different tuples for different perspectives. Also, we place the corresponding elements next to each marker token within $P_{i}$ as follows:
$y^{(2)}_{i}=\begin{aligned} &P^{(2)}_{i,1}\otimes tuple_{1}\,\text{[SSEP]}\,\ldots\,P^{(2)}_{i,K_{i}}\otimes tuple_{K_{i}},\end{aligned}$
where $P^{(2)}_{i,K_{i}}$ represents $K_{i}^{th}$ view in $P^{(2)}_{i}$ and $tuple_{K_{i}}$ is the $K_{i}^{th}$ sentiment tuple for given instance. $\otimes$ denotes an interleaved combination between marker tokens and elements. Detailed examples for both stages are present in Appendix E. We design the loss function for training the T5 model in second stage as follows.
@@ 2.3 Two-stage inference
During inference time, two stages are conducted sequentially. In the first stage, the model generates the initial order template, denoted as $\hat{y^{(1)}}$. In the second stage, we count the number of generated views from $\hat{y^{(1)}}$ to set $\hat{K}$. Using $\hat{K}$, we sample the top $\hat{K}$ views from the newly ranked set of views and constructs the final order template, referred to as $\hat{P}$. Finally, $\hat{P}$ is directly appended to the inference sentence, enabling the generation of different sentiment tuples for each view in $\hat{P}$. Throughout inference, we employ a constrained-decoding strategy De Cao et al. (2020) to ensure that the output at each stage conforms to the required format. The overall two-stage process is described in Figure 2.
@@ 3 Experiment
Methods ASQP ACOS MEMD Avg Time(s)
R15 R16 Lap Rest M-Rest M-Lap Books Clothing Hotel
TAS-BERT 34.78 43.71 27.31 33.53 - - - - - - -
Extract-Classify 36.42 43.77 35.80 44.61 - - - - - - -
One-ASQP (large) - - 41.56 60.69 - - - - - - -
Seq2Path - - 42.97 58.41 - - - - - - -
AugABSA 50.01 60.88 - - - - - - - - -
SCRAP 49.93 62.48 - - - - - - - - -
Paraphrase 46.93 57.93 43.51 61.16 57.38 35.07 39.30 43.00 68.79 50.34 40.63
DLO 48.18 59.79 43.64 59.99 57.07 35.56 42.63 43.35 70.27 51.16 260.74
MvP 51.04 60.39 43.92 61.54 58.12 35.25 42.57 43.94 69.06 51.76 2161.81
GPT-4o 40.45 47.29 24.77 46.53 35.11 20.69 30.39 40.27 24.84 34.48 -
LLaMa-3.1-8b 37.52 47.60 40.07 54.06 38.10 31.16 28.62 32.21 44.62 39.33 -
Qwen-2.5-7b 29.93 39.34 12.48 33.56 25.63 24.13 17.77 18.09 38.03 26.66 -
Mistral-7b 44.14 51.96 39.02 53.02 41.28 26.80 26.54 21.81 40.35 38.32 -
DOT (Ours) 51.91 61.24 44.92 59.25 58.25 39.02 43.02 43.37 69.94 52.28 298.17
Table 1: F1 scores for ABSA on nine datasets. Best results are in bold, second-best underlined. Results are averaged over five seeds. Time denotes average inference time.
@@ 3.1 Benchmark Datasets
We adopt two widely used ABSA datasets: ASQP and ACOS, where the task is to predict sentiment quadruples. For ASQP task, we use rest15 (R15) and rest16 (R16) datasets released from Pontiki et al. (2016); Zhang et al. (2022). For ACOS task, we use laptop16(Lap) and rest16(Rest) datasets constructed by Cai et al. (2021a); Pontiki et al. (2016). Also, we adopt additional ACOS benchmarks from MEMD datasets (Restaurant, Laptop, Books, Clothing, Hotel) Xu et al. (2023) which use a different source from the previous datasets. We refer to the Restaurant and Laptop datasets in MEMD as M-Rest and M-Laptop, respectively, for the sake of clarity.
@@ 3.2 Baselines
We benchmark our approach against a suite of extraction- and generation-based baselines. For a subset of baselines, we confine evaluation to the four datasets—R15, R16, Laptop, and Restaurant—because these methods were originally optimized for ASQP or ACOS tasks. Other baselines are assessed on their full applicable dataset range. This selective restriction ensures a fair comparison. The complete list of baselines is as follows: TAS-BERT Wan et al. (2020) jointly extracts and detects sentimental tuples. Extract-Classify Cai et al. (2021b) divide the task into two stages: extraction and classification. One-ASQP (large) Zhou et al. (2023) identify the aspect-opinion-sentiment (AOS) triplets simultaneously. Seq2Path Mao et al. (2022) generates sentiment tuples as multiple paths of a tree, and automatically selects a valid one. AugABSA Wang et al. (2023) generates a original text based on augmented sentiment quadruples. SCRAP Kim et al. (2024) optimizes the model to generate extract-then-assign reasonings and the corresponding sentiment quadruplets in sequence. Paraphrase Zhang et al. (2021a)  formulates a paraphrase generation process for ABSA with a single fixed order. DLO Hu et al. (2022)  augments data via the multiple order templates. MvP aggregates sentiment tuples generated from multiple orders of prompts via ensembling. Also, we benchmark popular LLMs such as GPT-4o Hurst et al. (2024), LLaMa-3.1-8b Dubey et al. (2024), Qwen-2.5-7b Yang et al. (2025), and Mistral-7b Jiang et al. (2023).
Detailed setups for LLMs are in Appendix H.
@@ 3.3 Implementation Details
We utilize the pre-trained T5-base (Raffel et al., 2020) model as the backbone for the first stage. We also use the model trained in the first stage as the backbone for the second stage, allowing us to leverage a tuned initial point for the ABSA dataset to have the regularization effect inspired by Fu et al. (2023).
Additionally, we observe that the label of the datasets (i.e. sentiment tuples) irregularly contains stop words. Thus, we eliminate irregularities in tuples through stop-word filtering in the second stage. We provide a detailed analysis and filtering process in Appendix A.
@@ 3.4 Results
Performance Comparison
We use the F1 score, which is a standard metric for ABSA, to measure the performance of the systems.
Table 1 demonstrates that our DOT framework achieves state-of-the-art results, ranking first on five of the nine benchmarks and second on three others. To further elaborate, a significant factor contributing to the enhanced performance is our approach of dividing the complex sentiment quadruple generation task into two subtasks, each handled by dedicated models. As a result, our method achieves better performance compared to a single model handling multi-view processing alone. Crucially, the effectiveness of this two-stage approach depends on the accuracy of the first task, which we comprehensively presented in Appendix C. However, its performance is slightly lower on the Rest and Clothing datasets, which we analyze in Section F.
@@
We also measure inference time using T5-base model for all baselines. We check inference time for each dataset, and average them.
As in Table 1, we dramatically reduce inference time particularly compared to the multi-view methods such as MvP Gou et al. (2023), by predicting solely the necessary number of views for each sample.
On the other hand, single view inferences Zhang et al. (2021a); Hu et al. (2022) and extraction approaches Wan et al. (2020); Cai et al. (2021b); Zhou et al. (2023) can be more memory- and time-efficientZhou et al. (2023), they generally exhibit inferior performance compared to generative methods. We provide more details on the inference time in Appendix D.
Transferability
@@
Table 2: Cross-dataset evaluation results for validating the effect of domain shift.
To examine the transferability of each model, we conduct an in-depth experiment on cross-dataset evaluation.
We group the datasets into two sources: SemEval Pontiki et al. (2016) (R15, R16, Rest) and Yelp (M-Rest), and assess performance by training on one group and testing on the other in a zero-shot setting. For the MvP model, we vary the number of views (3, 9, and 15) to evaluate sensitivity in static multi-view methods, while T5-paraphrase uses a static single order. As shown in Table 2, our model significantly outperforms the baselines in cross-dataset evaluation. Although T5-paraphrase suffers a smaller performance drop, it still falls behind our method, and MvP shows notable degradation regardless of view count. These results demonstrate that our model effectively identifies optimal views even for out-of-domain datasets.
Performance Across Model Scales
Beyond using T5-base as the sole backbone, we extend our experiments to examine how performance varies with model scale, specifically testing T5-small and T5-large. In this setup, we employ the same backbone model for both stages.
Figure 3: Bar chart illustrating F1 performance variations across T5-small, T5-base, and T5-large backbones on R15, R16, Laptop, and Rest benchmarks.
As shown in Figure 3, performance consistently improves across all benchmarks as the backbone scale increases, supporting the scaling law Kaplan et al. (2020) showing larger models yield better accuracy. Nonetheless, these gains do not fully compensate for the inefficiency introduced by loading larger models and performing multiple inference steps. Considering the trade-off between accuracy and computational cost, T5-base emerges as the most balanced choice.
Ablation study
To further investigate the effectiveness of each component of our framework, we conduct an ablation study and present the results in Table 3. Firstly, we record the results without stop-word filtering. Also, we unify the two stages into one, directly generating multiple order templates and tuples without including order prompting.
Additionally, we evaluate the results of sampling the views randomly, checking whether the entropy score is valid. Lastly, we exclude the multi-view approach by training and testing our model using only the view with the lowest entropy for each instance as the order template. We perform an ablation study by excluding one or more of the four components of our method mentioned earlier. By observing the gaps between these variants with the original model, we verify the effectiveness of each component of our method.
Model Configuration Average F1
Full Model 54.33
w/o filtering 53.31 (-1.02)
w/o stage division 52.73 (-1.60)
w/o entropy score 52.53 (-1.80)
w/o multi view 52.31 (-2.02)
w/o stage division, entropy score 50.04 (-4.29)
w/o filtering, stage division, entropy score 45.80 (-8.53)
Table 3: Ablation study for the proposed method, which shows the average F1 across ASQP and ACOS.